### v1

# Conduct Parameter Estimation in Homogeneous Goods Markets with Equilibrium Existence and Uniqueness Conditions: The Case of Log-linear Specification

Abstract
We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in a log-linear model with homogeneous goods markets. First, we derive such conditions. Second, Monte Carlo simulations confirm that in a log-linear model, incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters.
Keywords: Conduct parameters, Homogenous Goods Market, Mathematical Programming with Equilibrium Constraints, Monte Carlo simulation
JEL Codes: C5, C13, L1
1 Introduction
Measuring competitiveness is a crucial task in the empirical industrial organization literature. Conduct parameter is considered a useful measure of competitiveness. However, it cannot be directly measured from data because data usually lack information about marginal cost. Therefore, researchers aim to identify and estimate the conduct parameter.
As the simplest specification, Bresnahan (1982) considers identification of conduct parameters for the linear model. Matsumura and Otani (2023b, a) resolves the conflict on some identification problems between Bresnahan (1982) and Perloff and Shen (2012). However, researchers often implement alternative specifications, such as the log-linear model (Okazaki et al. 2022, Mérel 2009). In the context of the log-linear model, these papers identify estimation issues, with some estimated conduct parameters being unrealistically low or even negative. This raises doubts about the methodology, despite the identification strategy outlined by Lau (1982). This is an obstacle to choosing a better specification of the demand and supply functions.
To overcome the problem, we propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices as constraints. First, we prove that a unique equilibrium exists under certain theoretical conditions. As far as we know, this is a new result. Second, we show that incorporating equilibrium existence and uniqueness conditions resolves the above estimation problems and makes the model work in estimating conduct parameters even in the log-linear model. Hence, our results support those of Lau (1982) numerically.
2 Model
Consider data with $T$ markets with homogeneous products. Assume there are $N$ firms in each market. Let $t=1,\ldots,T$ be the index of markets. Then, we obtain the supply equation:
$P_{t}=-\theta\frac{\partial P_{t}(Q_{t})}{\partial Q_{t}}Q_{t}+MC_{t}(Q_{t}),$ (1)
where $Q_{t}$ is the aggregate quantity, $P_{t}(Q_{t})$ is the demand function, $MC_{t}(Q_{t})$ is the marginal cost function, and $\theta\in[0,1]$, which is the conduct parameter. The equation nests perfect competition $(\theta=0)$, Cournot competition $(\theta=1/N)$, and perfect collusion $(\theta=$ 1)1)1 ) (Bresnahan 1982).
Consider an econometric model. Assume that the demand and the marginal cost functions are written as follows:
$P_{t}=f(Q_{t},X^{d}_{t},\varepsilon^{d}_{t},\alpha),$
$MC_{t}=g(Q_{t},X^{c}_{t},\varepsilon^{c}_{t},\gamma),$
where $X^{d}_{t}$ and $X^{c}_{t}$ are the vector of exogenous variables, $\varepsilon^{d}_{t}$ and $\varepsilon^{c}_{t}$ are the error terms, and $\alpha$ and $\gamma$ are the vector of parameters. We also have the demand- and supply-side instruments, $Z^{d}_{t}$ and $Z^{c}_{t}$, and assume that the error terms satisfy the mean independence condition $E[\varepsilon^{d}_{t}\mid X^{d}_{t},Z^{d}_{t}]=E[\varepsilon^{c}_{t}\mid X^{c}_{t},Z^{c}_{t}]=0$.
2.1 Log-linear demand and log-linear marginal cost
Consider a log-linear model, which is a typical specification. The demand and marginal cost functions are specified as
$\log P_{t}$ $=\alpha_{0}-(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t},$ (2)
$\log MC_{t}$ $=\gamma_{0}+\gamma_{1}\log Q_{t}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t},$ (3)
where $Y_{t}$ is an excluded demand shifter, $W_{t}$ and $R_{t}$ are excluded cost shifters, and $Z_{t}^{R}$ is Bresnahan’s demand rotation instrument. Then, Equation (1) is written as
$P_{t}$ $=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+MC_{t}.$ (4)
By taking logarithm of Equation (4) and substituting Equation (3), we obtain
$\log P_{t}=-\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))+\gamma_{0}+\gamma_{1}\log Q_{t}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}.$ (5)
In this model, the number of equilibrium prices varies. Although it is a widely known model, to our knowledge, there is no paper examining this. Thus, we derive the conditions for the uniqueness in the following proposition which are used in our estimation.
Proposition 1.
Assume that $\alpha_{1}+\alpha_{2}Z^{R}_{t}\neq 0$. Let $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}$. The number of equilibrium prices $P_{t}^{*}>0$ is determined as follows:
  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, there is no equilibrium price,
  • When $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0,
    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\neq 1$, there is a unique equilibrium price,
    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, there are infinitely many equilibrium prices when $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, but there is no equilibrium price otherwise.
See the online appendix A.1 for the proof.
3 Estimation
To estimate parameters in the demand and supply equations, we use GMM estimation. Among GMM estimators, we apply the nonlinear system two-stage-least-squares (N2SLS) using Equation (2) and (5).
Let $\xi=(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\theta)$ be the vector of parameters. Given the demand equation (2) and the supply equation (5), we can write the error terms in the demand and supply equation as
${\varepsilon}_{t}^{d}(\xi)$ $=\log P_{t}-\alpha_{0}+(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}-\alpha_{3}\log Y_{t},$ (6)
${\varepsilon}_{t}^{c}(\xi)$ $=\log P_{t}+\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))-\gamma_{0}-\gamma_{1}\log Q_{t}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t}.$ (7)
To estimate the parameters, we convert the conditional moment conditions, $E[\varepsilon_{t}^{d}\mid Z_{t}^{d}]=E[\varepsilon_{t}^{c}\mid Z_{t}^{c}]=0$, into unconditional moment conditions, $E[\varepsilon_{t}^{d}Z_{t}^{d}]=E[\varepsilon_{t}^{c}Z_{t}^{c}]=0$. Using Equations (6) and (7), we construct the sample analog of the moment conditions:
$g(\xi)=\left[\begin{array}[]{l}\frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{d}_{t}(\xi)Z_{t}^{d}\\ \frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{c}_{t}(\xi)Z_{t}^{c}\end{array}\right].$
We define the GMM estimator as the vector $\xi^{*}$ that solves the problem,
$\min_{\xi}\ g(\xi)^{\top}Wg(\xi)$ (8)
where the weight matrix $W$ is defined as
$W=\left[\frac{1}{T}\sum_{t=1}^{T}Z_{t}^{\top}Z_{t}\right]^{-1}\text{ where }Z_{t}=\left[\begin{array}[]{ll}Z_{t}^{d\top}&0\\ 0&Z_{t}^{c\top}\end{array}\right].$ (11)
The solution to (8) is called the N2SLS estimator (Wooldridge 2010).
Using Proposition 1, we impose the following constraints:
$0\leq\theta\leq 1,$ (12)
$\alpha_{1}+\alpha_{2}Z_{t}^{R}>0,\quad\gamma_{1}>0,\quad t=1,\ldots,T$ (13)
$1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})>0,\quad t=1,\ldots,T.$ (14)
Constraint (12) is a standard assumption on the conduct parameter. Constraints (13) and (14) relate to the uniqueness of equilibrium prices. Constraint (13) implies the downward-sloping demand and upward-sloping marginal cost. See online appendix A.2 for the detailed setting.
4 Simulation results
We compare N2SLS estimation with and without Constraints (12), (13), and (14). Table 1 presents the results. First, as reported in the literature (Mérel 2009, Okazaki et al. 2022), N2SLS without constraints converges to implausibly low and negative conduct parameters. This is because the algorithm searches the area in which no equilibrium price and quantity exist and then finds unreasonable local optima that might have a lower objective function than at the true values. Second, N2SLS with constraints converges to the reasonable values. As the sample size increases, the bias and root-mean-square error (RMSE) decrease to levels comparable to linear models. Third, we find an ad hoc method that uses Equation (3) to compute the residual in the supply estimation, $\varepsilon_{t}^{c}$, and Equation (4) as an equality constraint with Constraints (12), (13), and (14). Then, the bias and RMSE of conduct parameter $\theta$ are reduced to 0.014 and 0.217 in Panel (c), although the results do not dominate the results in Panel (b) for all parameters. Therefore, incorporating equilibrium existence and uniqueness conditions simply is helpful for conduct parameter estimation in nonlinear models such as a log-linear model. See online appendix A.4 for additional experiments under different variances of errors and the results of a linear model.
Table 1: Performance comparison
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$1.070 7.012 -0.021 5.110 0.365 2.207 0.400 2.030
$\alpha_{1}$0.164 1.060 -0.001 0.782 0.073 0.458 0.096 0.574
$\alpha_{2}$0.011 0.104 -0.006 0.071 0.002 0.033 0.005 0.043
$\alpha_{3}$0.101 0.619 -0.005 0.474 0.021 0.198 0.029 0.187
$\gamma_{0}$ 9.735 15.743 9.636 10.870 13.173 13.269 13.294 13.351
$\gamma_{1}$0.070 1.624 -0.177 0.469 -0.184 0.248 -0.177 0.220
$\gamma_{2}$0.034 0.939 -0.098 0.317 -0.090 0.152 -0.080 0.127
$\gamma_{3}$0.047 0.750 -0.091 0.311 -0.098 0.156 -0.085 0.133
$\theta$3e+05 3e+06 -2e+05 2e+06 -8e+04 9e+04 -9e+04 1e+05
Runs converged (%) 99.500 99.800 98.600 98.400
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.905 6.954 0.120 5.001 0.072 2.042 0.052 1.563
$\alpha_{1}$0.141 1.053 0.018 0.768 0.010 0.313 0.008 0.241
$\alpha_{2}$0.006 0.101 0.000 0.062 0.001 0.024 0.001 0.019
$\alpha_{3}$0.088 0.620 0.007 0.475 -0.001 0.186 0.003 0.152
$\gamma_{0}$1.748 14.206 -0.938 6.428 0.015 1.995 0.163 1.570
$\gamma_{1}$ 0.254 1.927 0.129 0.825 0.018 0.226 0.003 0.170
$\gamma_{2}$ 0.117 1.083 0.049 0.467 0.008 0.148 0.007 0.116
$\gamma_{3}$ 0.098 0.890 0.058 0.478 -0.001 0.148 0.003 0.118
$\theta$0.100 0.441 -0.072 0.424 -0.121 0.351 -0.148 0.333
Runs converged (%) 99.600 99.900 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Note: The error terms are drawn from a normal distribution, $N(0,\sigma)$. True values: $\alpha_{0}=20.0,\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=1.0,\gamma_{0}=5.0,\gamma_{1}=1.0,\gamma_{2}=1.0,\gamma_{3}=1.0,\theta=0.5$ and $\sigma=1.0$. See online appendix A.2 for the setting.
5 Conclusion
We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in homogeneous goods markets. First, we derive the conditions. Second, Monte Carlo simulations confirm that incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters in a log-linear model, as reported in the literature.
Acknowledgments
We thank Jeremy Fox, Yelda Gungor, and Isabelle Perrigne for valuable comments. This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Appendix A Online appendix
A.1 Existence and uniqueness of equilibrium prices
Proposition 1 proposes the conditions for the unique existence of $P_{t}(>0)$ solving the demand equation (2) and supply equation (5) for $P_{t}$ under $\theta\in[0,1]$. The proof is not based on the optimization by individual firms but checks if there exists a point at which the demand function and the marginal cost function cross. Therefore, we allow an equilibrium price to exist when the demand and the marginal cost are upward-sloping.
Proof.
Rewriting the demand equation (2) as
$\log Q_{t}(P_{t})=\frac{\alpha_{0}-\log P_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{(\alpha_{1}+\alpha_{2}Z^{R}_{t})}$
and substituting this into the supply equation (5), we obtain
$P_{t}$ $=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp\left(\gamma_{0}+\gamma_{1}\log Q_{t}(P_{t})+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}\right).$
$=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp\left(\gamma_{0}+\gamma_{1}\frac{\alpha_{0}-\log P_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{(\alpha_{1}+\alpha_{2}Z^{R}_{t})}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}\right)$
$=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp\left(\Xi+\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\log P_{t}\right)$
$=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ (15)
where $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}$.
Any price that satisfies (15) becomes an equilibrium price. To find an equilibrium price $P_{t}^{*}$, we define $\Delta(P_{t})$ as follows:
$\Delta(P_{t})$ $=\underbrace{[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}}_{(I)}-\underbrace{\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}}_{(II)}.$ (16)
We label the first term as (I) and the second term as (II).
An important note is that $P_{t}=0$ satisfies $\Delta(P_{t})=0$. Thus $P_{t}^{*}=0$ holds. However, our interest is a unique positive equilibrium price, so we seek conditions in which there is a unique positive price that satisfies $\Delta(P_{t})>0$.
When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) is always negative on $P_{t}>0$. In contrast, (II) is non-negative regardless of the sign of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ on $P_{t}>0$. Therefore, $\Delta(P_{t})$ is always negative on $P_{t}>0$, which implies that there is no positive equilibrium price.
When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) becomes a line passing through the origin with a positive slope, illustrated in Figure 1. Note that the first and second derivatives of (II) are given as
$\frac{d}{dP_{t}}\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ $=-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\exp(\Xi)P_{t}^{-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1},$ (17)
$\frac{d^{2}}{dP_{t}^{2}}\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ $=-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\left(-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1\right)\exp(\Xi)P_{t}^{-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-2}.$ (18)
Therefore, the shape of (II) on $P_{t}>0$ is changed based on the value of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ and the number that (I) and (II) cross on $P_{t}>0$ is also determined.
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})<0$ is illustrated in Figure 1(a). In this case, (II) is a monotone decreasing convex function because the first and second derivatives are both negative. As (I) is monotone increasing in $P_{t}>0$, (I) and (II) cross only once on $P_{t}>0$.
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\in[0,1)$ is illustrated in Figure 1(b). In this case, (II) becomes a monotone increasing concave function passing through the origin because the first derivative is positive and the second derivative is negative. As the value of the first derivative is infinity at $P_{t}=0$ since $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})-1<0$, (II) should be greater than (I) around $P_{t}=0$. Then as $P_{t}$ becomes large starting from zero, the difference in both terms becomes small, and both terms cross only once on $P_{t}>0$ eventually.
When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, (16) becomes
$0=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)]P_{t}.$
If $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})=\exp(\Xi)$, the above equation holds for any $P_{t}$. Thus there are infinitely many equilibrium prices on $P_{t}>0$, which is illustrated in Figure 1(c). When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq\exp(\Xi)$, there is no equilibrium price on $P_{t}>0$ because the right-hand side of the equation is always non-zero on $P_{t}>0$, which is illustrated in Figure 1(d).
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})>1$ is illustrated in Figure 1(e). In this case, (II) becomes a monotone increasing convex function passing through the origin because the first and second derivatives are positive. Therefore, (I) and (II) cross only once on $P_{t}>0$.
Figure 1: The illustrations of the equilibrium existence and uniqueness conditions
$P_{t}$$\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$$[1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})]P_{t}$
(a) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}<0$
$P_{t}$
(b) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\in[0,1)$
$P_{t}$
(c) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}=1,1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})=\exp(\Xi)$
$P_{t}$
(d) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}=1,1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})\neq\exp(\Xi)$
$P_{t}$
(e) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}>1$
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. The thick line represents the first term and the dotted line represents the second term in (16).
From $\Delta(P_{t})=0$, the positive equilibrium price can be written as
00 0 $=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$
00 0 $=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$
$P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$ $=\frac{1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})}{\exp(\Xi)}$
$P_{t}^{*}$ $=\left(\frac{1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})}{\exp(\Xi)}\right)^{-\frac{\alpha_{1}+\alpha_{2}Z^{R}_{t}}{\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}_{t}}}.$ (19)
Figure 2 illustrates how the demand and supply equations cross under the conditions. When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the supply equation (5) is ill-defined because the inside of the log function becomes negative. Thus, there should not be any equilibrium. Hereafter, assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand equation and supply equation coincides. Thus there are infinitely many equilibria in the model. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand and supply equations have a same slope and different intercepts, which means that both equations become parallel. Thus, there is no equilibrium. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, the demand and supply equations have different slopes, and hence we can find a unique equilibrium.
Figure 2: The illustrations of the equilibrium existence and uniqueness conditions
[…]
(a) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
[…]
(b) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
[…]
(c) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$, and $\gamma_{1}>0$
[…]
(d) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}\leq 0$, and $\gamma_{1}>0$
[…]
(e) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$, and $\gamma_{1}\leq 0$
[…]
(f) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}\leq 0$, and $\gamma_{1}\leq 0$
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. Case (a) and (b) also allow both the demand and supply equations to be downward sloping.
A.2 Simulation and estimation procedure
To generate the simulation data, for each model, we first generate the exogenous variables $Y_{t},Z^{R}_{t},W_{t},R_{t},H_{t}$, and $K_{t}$ and the error terms $\varepsilon_{t}^{c}$ and $\varepsilon_{t}^{d}$ based on the data generation process in Table 2. By substituting the Equation (2) into Equation (5) and solving it for $P_{t}$, the log aggregate quantity is given as:
$\log Q_{t}$ $=\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))-\gamma_{0}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t}+\varepsilon^{d}_{t}-\varepsilon^{c}_{t}}{\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}_{t}}.$ (20)
We compute the equilibrium quantity $Q_{t}$ for the log-linear model by (20). We then compute the equilibrium price $P_{t}$ by substituting $Q_{t}$ and other variables into the demand function (2). We generate 1000 data sets of 100, 200, 1000, 1500 markets. We jointly estimate the demand and supply parameters by the simultaneous equation model (Wooldridge 2010) from the true values. We use state-of-the-art constrained optimization solvers, i.e., Ipopt.jl which implements an interior point line search filter method that aims to find a local solution of nonlinear programming problems.
Table 2: True parameters and distributions
$\alpha_{0}$ $20.0$
$\alpha_{1}$ $1.0$
$\alpha_{2}$ $0.1$
$\alpha_{3}$ $1.0$
$\gamma_{0}$ $5.0$
$\gamma_{1}$ $1.0$
$\gamma_{2}$ $1.0$
$\gamma_{3}$ $1.0$
$\theta$ $0.5$
Demand shifter
$Y_{t}$ $N(0,1)$
Demand rotation instrument
$Z^{R}_{t}$ $U(0,1)$
Cost shifter
$W_{t}$ $U(1,3)$
$R_{t}$ $U(1,3)$
$H_{t}$ $W_{t}+U(0,1)$
$K_{t}$ $R_{t}+U(0,1)$
Error
$\varepsilon^{d}_{t}$ $N(0,\sigma)$
$\varepsilon^{c}_{t}$ $N(0,\sigma)$
Note: $\sigma=\{0.5,1.0,2.0\}$. $N:$ Normal distribution. $U:$ Uniform distribution.
A.3 Why does the estimation without equilibrium conditions converge to the extremely low conduct parameters?
We should interpret reasons why the estimation without equilibrium conditions converges to the negative conduct parameters, as shown in Panel (a) in Table 1, although the obtained parameters cannot be interpreted as the exogenous elements determining equilibrium outcomes. First, we rewrite the terms in Equation (7) into
$\log(1-\theta C_{t})-\gamma_{0}$
where $C_{t}=\alpha_{1}+\alpha_{2}Z^{R}_{t}$ is given in the demand estimation. If the model allows $\theta<0$, then $1-\theta C_{t}$ can take an arbitrarily large value through log-transformation relative to $\gamma_{0}$. For example, if $0001-\theta C_{t}=1,000,$ 000, then $\log(1-\theta C_{t})=13.8$, which has a small contribution relative to $\gamma_{0}$ for evaluating the moment conditions.
A.4 Additional experiments
Additional results for different $\sigma$ are shown in Tables 3 and 4. As a summary, the main findings in the main text are robust.
In Tables 6 and 7, we illustrate that N2SLS with Constraints (12) works for the linear model as in Matsumura and Otani (2023b). This means that incorporating the conditions is innocuous for the linear model.
Additional ad hoc improvement is possible by using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14). Table 5 shows that the estimation of the conduct parameter $\theta$ improves, that is, the bias is 0.014 and RMSE is 0.217. However, the results with ad hoc improvement do not dominate the results without the improvement for all parameters.
Table 3: Performance comparison ($\sigma=0.5$)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$ 0.119 4.055 0.346 3.180 0.344 1.497 0.379 1.582
$\alpha_{1}$ 0.019 0.620 0.054 0.486 0.076 0.429 0.099 0.576
$\alpha_{2}$ 0.000 0.050 0.003 0.039 0.002 0.029 0.003 0.036
$\alpha_{3}$ 0.009 0.336 0.025 0.268 0.028 0.129 0.031 0.129
$\gamma_{0}$ 10.104 11.066 9.397 10.288 13.240 13.280 13.314 13.346
$\gamma_{1}$0.146 0.340 -0.166 0.264 -0.180 0.198 -0.181 0.192
$\gamma_{2}$0.072 0.229 -0.079 0.165 -0.087 0.108 -0.086 0.099
$\gamma_{3}$0.069 0.230 -0.082 0.171 -0.088 0.110 -0.087 0.100
$\theta$5e+04 1e+05 -3e+04 6e+04 -8e+04 9e+04 -8e+04 9e+04
Runs converged (%) 100.000 100.000 98.100 98.700
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.006 3.556 0.175 3.158 0.059 0.977 -0.017 0.763
$\alpha_{1}$0.002 0.545 0.026 0.484 0.009 0.150 -0.003 0.117
$\alpha_{2}$ 0.002 0.044 0.004 0.037 0.000 0.012 0.000 0.010
$\alpha_{3}$ 0.001 0.309 0.012 0.266 0.005 0.091 0.000 0.073
$\gamma_{0}$0.616 3.317 -0.221 2.191 0.275 1.076 0.421 0.972
$\gamma_{1}$ 0.089 0.399 0.038 0.252 -0.015 0.107 -0.027 0.089
$\gamma_{2}$ 0.044 0.265 0.022 0.170 -0.007 0.074 -0.011 0.061
$\gamma_{3}$ 0.046 0.273 0.019 0.178 -0.008 0.078 -0.012 0.059
$\theta$0.064 0.390 -0.079 0.339 -0.119 0.260 -0.145 0.261
Runs converged (%) 99.900 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Table 4: Performance comparison ($\sigma=2.0$)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$2.358 11.286 -1.457 8.696 0.231 4.747 0.450 4.028
$\alpha_{1}$0.360 1.770 -0.227 1.313 0.040 0.744 0.095 0.711
$\alpha_{2}$0.021 0.138 -0.012 0.142 0.002 0.063 0.006 0.065
$\alpha_{3}$0.175 0.967 -0.138 0.867 0.006 0.434 0.031 0.347
$\gamma_{0}$ 9.868 21.289 10.490 13.359 13.151 13.453 13.177 13.384
$\gamma_{1}$0.015 2.542 -0.130 1.042 -0.202 0.400 -0.186 0.339
$\gamma_{2}$0.057 1.656 -0.083 0.735 -0.090 0.274 -0.099 0.224
$\gamma_{3}$0.020 2.115 -0.065 0.719 -0.097 0.279 -0.092 0.232
$\theta$7e+05 5e+06 -3e+05 2e+06 -1e+05 4e+05 -9e+04 1e+05
Runs converged (%) 99.600 99.800 98.300 98.900
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$2.008 11.292 -1.149 8.680 0.145 4.655 0.274 3.838
$\alpha_{1}$0.308 1.773 -0.181 1.313 0.020 0.713 0.041 0.590
$\alpha_{2}$0.016 0.136 -0.007 0.139 0.003 0.057 0.004 0.043
$\alpha_{3}$0.147 0.985 -0.113 0.875 -0.003 0.429 0.017 0.333
$\gamma_{0}$ 0.661 11.910 -1.090 13.475 -0.665 4.501 -0.404 3.245
$\gamma_{1}$0.044 1.600 0.172 1.770 0.098 0.556 0.066 0.391
$\gamma_{2}$0.056 1.086 0.088 1.172 0.054 0.346 0.024 0.252
$\gamma_{3}$0.020 1.205 0.108 1.225 0.049 0.365 0.031 0.267
$\theta$0.229 0.475 -0.163 0.464 -0.090 0.410 -0.108 0.394
Runs converged (%) 99.200 99.200 99.700 100.000
Sample size ($T$) 100 200 1000 1500
Table 5: Ad hoc method using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.347 2.978 -0.153 2.547 0.065 0.976 -0.009 0.760
$\alpha_{1}$0.043 0.452 -0.016 0.388 0.010 0.150 -0.001 0.117
$\alpha_{2}$0.024 0.065 -0.017 0.054 0.000 0.012 0.000 0.010
$\alpha_{3}$0.028 0.265 -0.013 0.224 0.005 0.091 0.001 0.073
$\gamma_{0}$2.458 5.660 -1.258 2.959 -0.132 0.942 -0.153 0.748
$\gamma_{1}$ 0.112 0.602 0.057 0.260 0.014 0.103 0.014 0.082
$\gamma_{2}$ 0.053 0.314 0.030 0.172 0.007 0.073 0.010 0.060
$\gamma_{3}$ 0.061 0.426 0.028 0.181 0.006 0.077 0.008 0.058
$\theta$ 0.169 0.407 0.109 0.356 -0.003 0.155 0.017 0.116
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.835 9.908 -0.561 6.310 0.147 4.424 0.301 3.686
$\alpha_{1}$0.109 1.563 -0.076 0.935 0.024 0.677 0.047 0.566
$\alpha_{2}$0.043 0.124 -0.033 0.135 -0.005 0.061 0.001 0.045
$\alpha_{3}$0.047 0.909 -0.078 0.664 -0.004 0.404 0.018 0.323
$\gamma_{0}$6.629 20.418 -5.373 16.692 -1.445 4.738 -1.123 3.411
$\gamma_{1}$ 0.422 2.630 0.384 2.103 0.137 0.580 0.115 0.409
$\gamma_{2}$ 0.184 1.788 0.185 1.318 0.073 0.357 0.048 0.261
$\gamma_{3}$ 0.199 1.885 0.215 1.496 0.068 0.376 0.055 0.276
$\theta$ 0.225 0.459 0.227 0.447 0.098 0.357 0.084 0.313
Runs converged (%) 99.200 99.800 99.800 100.000
Sample size ($T$) 100 200 1000 1500
Table 6: N2SLS without Constraints (12) Results of the linear model
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.018 0.465 0.007 0.323 -0.008 0.213 -0.006 0.097
$\alpha_{1}$0.045 2.257 0.024 1.523 0.018 1.016 -0.031 0.455
$\alpha_{2}$0.001 0.255 -0.001 0.176 -0.004 0.115 0.001 0.051
$\alpha_{3}$0.005 0.108 0.003 0.075 -0.001 0.050 -0.001 0.022
$\gamma_{0}$0.061 0.732 -0.005 0.474 -0.021 0.346 -0.005 0.152
$\gamma_{1}$0.311 3.450 -0.124 1.928 -0.081 1.303 -0.003 0.548
$\gamma_{2}$ 0.009 0.109 -0.001 0.071 0.003 0.051 0.000 0.023
$\gamma_{3}$ 0.001 0.108 0.003 0.075 0.003 0.053 0.000 0.022
$\theta$ 0.047 0.354 0.017 0.209 0.014 0.135 0.003 0.058
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.027 1.023 -0.002 0.641 -0.004 0.448 -0.016 0.188
$\alpha_{1}$0.024 4.396 -0.169 2.965 0.061 2.060 0.011 0.905
$\alpha_{2}$0.006 0.494 0.016 0.325 -0.007 0.234 -0.006 0.100
$\alpha_{3}$0.006 0.223 0.002 0.153 -0.001 0.099 -0.003 0.045
$\gamma_{0}$0.318 1.769 -0.091 1.059 -0.086 0.714 -0.004 0.308
$\gamma_{1}$ 5.859 210.853 -0.679 6.280 -0.338 2.972 -0.050 1.110
$\gamma_{2}$ 0.035 0.247 0.011 0.157 0.011 0.104 0.000 0.045
$\gamma_{3}$ 0.045 0.250 0.012 0.150 0.010 0.104 0.002 0.045
$\theta$0.399 18.450 0.098 0.738 0.054 0.308 0.009 0.113
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.263 2.596 0.071 1.670 -0.040 0.947 -0.002 0.412
$\alpha_{1}$0.271 10.820 0.008 6.492 0.236 4.263 0.021 1.809
$\alpha_{2}$0.044 1.253 0.023 0.779 -0.031 0.483 -0.003 0.210
$\alpha_{3}$0.024 0.584 0.008 0.343 -0.004 0.225 0.003 0.092
$\gamma_{0}$2.074 19.624 -0.551 3.043 -0.171 1.516 -0.051 0.633
$\gamma_{1}$ 58.209 1750.688 -2.416 56.909 -3.617 39.044 -0.103 2.334
$\gamma_{2}$ 0.242 2.430 0.065 0.409 0.020 0.220 0.006 0.093
$\gamma_{3}$ 0.230 2.328 0.055 0.404 0.010 0.219 0.008 0.092
$\theta$6.668 233.851 0.372 6.334 0.418 3.820 0.024 0.245
Sample size ($T$) 50 100 200 1000
Note: The data generating process follows Matsumura and Otani (2023b).
Table 7: N2SLS with Constraints (12) for the linear model
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.013 0.462 0.008 0.322 -0.008 0.213 -0.006 0.097
$\alpha_{1}$0.096 2.201 0.015 1.511 0.018 1.016 -0.031 0.455
$\alpha_{2}$ 0.006 0.247 0.001 0.174 -0.004 0.115 0.001 0.051
$\alpha_{3}$0.004 0.108 0.003 0.074 -0.001 0.050 -0.001 0.022
$\gamma_{0}$0.054 0.724 -0.002 0.472 -0.021 0.346 -0.005 0.152
$\gamma_{1}$0.098 2.620 -0.093 1.847 -0.081 1.303 -0.003 0.548
$\gamma_{2}$ 0.008 0.108 -0.002 0.070 0.003 0.051 0.000 0.023
$\gamma_{3}$ 0.001 0.107 0.003 0.075 0.003 0.053 0.000 0.022
$\theta$ 0.023 0.258 0.014 0.197 0.014 0.135 0.003 0.058
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$ 0.012 1.015 0.012 0.636 0.001 0.446 -0.016 0.188
$\alpha_{1}$0.383 4.216 -0.291 2.872 0.015 2.012 0.011 0.905
$\alpha_{2}$ 0.042 0.466 0.033 0.310 -0.001 0.227 -0.006 0.100
$\alpha_{3}$ 0.000 0.222 0.004 0.152 0.000 0.099 -0.003 0.045
$\gamma_{0}$0.233 1.612 -0.067 1.020 -0.076 0.703 -0.004 0.308
$\gamma_{1}$ 0.406 4.332 -0.024 3.174 -0.178 2.585 -0.050 1.110
$\gamma_{2}$ 0.025 0.231 0.008 0.154 0.010 0.103 0.000 0.045
$\gamma_{3}$ 0.034 0.233 0.010 0.146 0.009 0.103 0.002 0.045
$\theta$ 0.014 0.380 0.018 0.311 0.035 0.260 0.009 0.113
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.138 2.592 0.142 1.671 0.004 0.936 0.000 0.410
$\alpha_{1}$0.986 11.090 -0.665 6.325 -0.174 4.042 0.003 1.788
$\alpha_{2}$ 0.065 1.255 0.112 0.756 0.023 0.447 0.000 0.207
$\alpha_{3}$0.006 0.589 0.019 0.345 0.003 0.224 0.003 0.092
$\gamma_{0}$0.123 3.350 -0.298 2.455 -0.081 1.379 -0.047 0.628
$\gamma_{1}$ 0.635 7.424 0.141 5.668 -0.081 4.255 -0.040 2.159
$\gamma_{2}$ 0.016 0.484 0.037 0.349 0.009 0.205 0.005 0.093
$\gamma_{3}$ 0.009 0.517 0.026 0.336 0.000 0.208 0.007 0.092
$\theta$0.029 0.446 0.025 0.414 0.034 0.381 0.016 0.220
Sample size ($T$) 50 100 200 1000
Note: The data generating process follows Matsumura and Otani (2023b).